# The Landscape of Collapsar Outflows: Structure, Signatures and Origins of Einstein Probe Relativistic Supernova Transients

### v1

@@ 1 Introduction
One natural candidate for producing the sub- and mildly relativistic hot plasma is the “cocoon,” generated during the jet’s propagation within the collapsing star (Mészáros & Rees, 2001; Ramirez-Ruiz et al., 2002; Matzner, 2003; Lazzati & Begelman, 2005). Cocoon emission has been predicted to produce vigorous UV/optical signals (Nakar & Piran, 2017), potentially associated with SNe Ib/c (Piran et al., 2019), and is hypothesized as a possible source of fast optical transients (Gottlieb et al., 2022c). Indeed, cocoon emission is a popular interpretation for several puzzling EP transients, including EP240414a (Hamidani et al., 2025b; van Dalen et al., 2025; Zheng et al., 2025), EP240801a (Jiang et al., 2025), EP241021a (Gianfagna et al., 2025), and EP250108a (Eyles-Ferris et al., 2025; Yadav et al., 2025) [^1] Although Yadav et al. (2025) argue against quasi-spherical mildly relativistic cocoons, we note that mildly relativistic cocoons must inherently exhibit some degree of collimation. In fact, the cocoon structure naturally arising during jet propagation constitutes the “structured jet” proposed by Jiang et al. (2025); Yadav et al. (2025) (see Gottlieb et al., 2021, for further discussion on the cocoon-jet interplay)..
@@ 2 Physics of Collapsars
The collapse of rapidly rotating stars leads to the formation of a spinning PM (e.g., Raynaud et al., 2020; Masada et al., 2022; White et al., 2022), which may eventually collapse into a spinning BH. Such a BH can launch relativistic jets via the Blandford-Znajek (BZ; Blandford & Znajek, 1977) mechanism. Sustained jet launching capable of powering GRBs through the BZ mechanism (Popham et al., 1999; MacFadyen & Woosley, 1999) typically requires three conditions (e.g., Symbalisty, 1984; Metzger et al., 2008; Gottlieb et al., 2022a): (i) a dimensionless BH spin parameter $a\gtrsim 0.3$ (Gottlieb et al., 2023); (ii) the presence of a steady accretion disk surrounding the BH; and (iii) sufficient poloidal magnetic flux available to the BH. Gottlieb et al. (2024) demonstrated that an accretion disk may form around the PM before its collapse; such a disk can inhibit magnetic reconnection, enabling the BH to inherit the PM’s magnetic flux and thus relaxing the third requirement. Consequently, an accretion disk around a moderately spinning BH emerges as the only prerequisite for BZ jet launching. However, the angular momentum profiles of stellar progenitors and the likelihood of disk formation during the PM phase remain uncertain (see e.g., Fryer et al., 2025).
@@
Structure: The collapse results in the formation of a proto-magnetar, which launches subrelativistic outflows (violet) prior to collapsing into a black hole. The ensuing formation of a black hole surrounded by an accretion disk drives moderately collimated, magnetized disk ejecta (turquoise) and a wobbling ultra-relativistic episodic jet powered by the Blandford-Znajek mechanism (yellow). As the jet propagates, it shocks the surrounding stellar material, generating a collimated, mildly relativistic cocoon (blue). An additional subrelativistic supernova shockwave from core bounce (magenta) is possible; however, the proto-magnetar and/or disk ejecta are consistent with the energy and velocity distributions observed in stripped-envelope supernovae, rendering the supernova shockwave unnecessary to account for observations. We note that the illustrated collimation of the cocoon and disk ejecta represents characteristic angles where the bulk of the energy is concentrated. In practice, the outflows mix in a spherical morphology, with the energy distributed as a function of angle.
@@ 2.1.3 Wobbling Jets and Cocoon
If the BH possesses a moderate spin of $a\gtrsim 0.2$, a sufficient amount of magnetic flux will activate the BZ mechanism, launching bipolar relativistic jets. The jets interact with the dense stellar core, producing a shocked cocoon composed of mildly relativistic jet plasma and spherical non-relativistic shocked stellar material (Bromberg et al., 2011). As the cocoon expands, it envelops and interacts with the accretion disk. Large-scale, long-duration, high-resolution 3D general-relativistic magnetohydrodynamic collapsar simulations by Gottlieb et al. (2022b) have demonstrated the self-consistent formation of both the accretion disk and jets from the collapsing star and through the BZ mechanism. These simulations revealed that jet–disk interactions transfer stochastic momentum to the disk, destabilizing it and continuously altering its orientation. Consequently, the jet, which is anchored to the disk’s angular momentum, develops a wobbling motion at its base. Figure 2, taken from Gottlieb et al. (2022b)’s 3D simulations, demonstrates that this motion is imprinted onto the jet’s larger-scale structure and remains observable even after the jet emerges from the collapsing star on a breakout time of a few seconds.
@@ 2.2 Outflow Structure
The relative contribution between the jet-cocoon structure and disk ejecta depends on the BH spin. Provided sufficient magnetic flux is supplied to the BH, either by the PM (Gottlieb et al., 2024) or through dynamo processes (Shibata et al., 2025), two regimes are relevant: [^2] The jet power depends on the BH spin and the magnetic flux. Since collapsar jet launching requires a high magnetic flux (Issa et al., 2025), this imposes a constraint on the BH spin, as high spins would produce jets with luminosities exceeding those observed in GRBs. As a result, spins of $a\gtrsim 0.5$ are ruled out (Gottlieb et al., 2023). moderately spinning BHs ($0.2\lesssim a\lesssim 0.5$) that launch GRB jets; and slowly spinning BHs ($a\lesssim 0.2$) that do not produce jets capable of powering GRBs. At $a\approx 0.25$, the power of the disk ejecta is comparable to that of the BZ jets, $L_{\rm disk}\approx L_{\rm BZ}$ (McKinney et al., 2012; Narayan et al., 2012; Tchekhovskoy & McKinney, 2012; Bopp & Gottlieb, 2025). Given that jets require a persistent accretion disk, the disk and jet launching durations are also expected to be comparable, leading to similar total energies, $E_{\rm disk}\approx E_{\rm BZ}$. Since jets deposit about half of their energy into the cocoon (Bromberg et al., 2011; Gottlieb et al., 2021), we expect $E_{\rm disk}\approx 2E_{\rm cocoon}\approx 2E_{\rm jet}$. Conversely, in the case of slowly spinning BHs, the disk ejecta becomes the dominant energy channel, $E_{\rm disk}\gg E_{\rm jet}$.
@@
Ultimately, all outflow components interact and merge into a single structured outflow with both radial and angular profiles. Figure 3(a) presents the radially integrated isotropic-equivalent energy as a function of angular separation from the stellar rotation axis. The resulting global angular structure features a flat core with power-law wings (Gottlieb et al., 2021), and remains unaffected by jet wobbling (Gottlieb et al., 2022b), as it reflects radial and azimuthal integrations over multiple jet episodes launched at varying angles. However, the angular structure of individual jet episodes has not been investigated until now. Figure 3(b,c) depicts $E_{\rm iso}$ (b) and Lorentz factor (c) across a shell containing a single bipolar wobbling jet episode. Both jets show a prominent deviation from the rotational axis due to the wobbling motion. In stark contrast to the smooth core+power-law structure seen in the global outflow structure (dashed lines in panel b), single jet episodes exhibit a distinct core and steep power-law wings, whose power-law indices exceed $\sim 5$ for both the energy and Lorentz factor. Accordingly, the angular profile of single jet episodes is more characteristic of a top-hat jet than a structured jet. We note that the relatively low jet Lorentz factor of $\Gamma\lesssim 10$ in Fig. 3(c) arises from the floor values imposed in numerical simulations, which constrain the asymptotic Lorentz factor. A higher Lorentz factor, as suggested by observations, is expected to yield an even more top-hat-like structure.
@@ 3.2.1 Challenges of Proposed Models
The detection of GRBs, along with the direct association of some events with SNe Ic-BL, supports a collapsar origin and thus motivates a focus on collapsar models rather than other events such as accreting white dwarfs or binaries (Wu et al., 2025), or tidal disruption events (Xinwen et al., 2025), whose afterglows evolve over considerably longer timescales (e.g., Alexander et al., 2020). As density fluctuations in the CSM cannot account for the observed sharp features (Nakar & Granot, 2007), a few alternative models have been proposed to explain the afterglow rebrightening. Two widely discussed models proposed to explain the rebrightening are off-axis emission from a jet with angular structure (e.g., Yadav et al., 2025; Zheng et al., 2025), and refreshed shocks (e.g., Busmann et al., 2025; Ricci et al., 2025; Srivastav et al., 2025), or a combination of both (Gianfagna et al., 2025). However, both of these models face significant challenges.
Structured jets can naturally explain the absence of $\gamma$-ray emission when the observer’s line of sight initially misses the jet core, and the rebrightening emerges as the jet core gradually enters the line of sight. However, the very steep rise observed in EP241021a ($L_{\nu}\sim t^{4}$; Busmann et al., 2025) and EP240414a ($L_{\nu}\sim t^{3}$; Zheng et al., 2025), cannot be reproduced by a smooth angular jet structure (Lamb et al., 2021), unless the jet is composed of discrete segments (Gianfagna et al., 2025) with distinct Lorentz factors (Xinwen et al., 2025), similar to top-hat jets (Granot et al., 2002; Nakar et al., 2002). Such a segmented configuration of two-component relativistic jets (see Granot, 2005; Beniamini et al., 2020; Li et al., 2023), previously proposed to explain rebrightening features (Liang et al., 2013), naturally arises from wobbling top-hat jets, as we demonstrated in §2.2, and as we argue in §3.2.2, likely accounts for the observed sharp rebrightening behavior. In contrast, the assumed structures in Gianfagna et al. (2025); Yadav et al. (2025); Zheng et al. (2025) of jets with only an angular profile formed through continuous interaction with the ambient medium, will produce a monotonic jet+cocoon structure (see jet structure profiles in Gottlieb et al., 2021), leading to a gradual and prolonged rise, similar to that in GW170817 (e.g., Mooley et al., 2018). Similar challenges arise for radial structures, which are characteristic of choked jets (see §3.3.2).
@@ 3.2.2 Wobbling Jets
The high-redshift ($z=4.9$) event EP240315a exhibited a rebrightening at $t_{\rm rb}\approx 13\,{\rm hr}$. Assuming the kinetic energy is comparable to the high-energy emission, $E_{k,\rm iso}\approx E_{\rm iso}\approx 10^{54}\,\,{{\rm erg}}$, and $n=1\,\,{{\rm cm}}^{-3}$, we find from Equation (1) that $\Gamma\approx 20$, implying the observer was located $\Delta\theta\approx 0.05\,{\rm rad}$ off-axis from the most energetic segment of the wobbling jet. The jet afterglow in EP240414a at $z=0.4$ indicates $E_{k,\rm iso}\approx 10^{51}\,\,{{\rm erg}}$ with $n=1\,\,{{\rm cm}}^{-3}$ for the initially observed decelerating jet prior to brightening (Sun et al., 2025). Assuming a typical total (over all episodes) jet isotropic equivalent energy of $E_{k,\rm iso}\approx 10^{53}\,\,{{\rm erg}}$ at $t_{\rm rb}\approx 1\,{\rm day}$, we find $\Gamma\approx 7$, implying that $\Delta\theta\approx 0.15\,{\rm rad}$. For EP241021a ($z=0.75$), which exhibited a rebrightening at $t_{\rm rb}=6\,{\rm days}$, we adopt $E_{k,\rm iso}=10^{53}\,\,{{\rm erg}}$ and $n=10^{-3}\,\,{{\rm cm}}^{-3}$ based on afterglow modeling by Yadav et al. (2025). This yields $\Gamma\approx 9$, corresponding to $\Delta\theta\approx 0.1\,{\rm rad}$. A secondary, weaker rebrightening, which could be powered by a weaker wobbling jet episode, was observed at $t_{\rm rb}=20.7\,{\rm days}$ (Xinwen et al., 2025), which corresponds to $\Delta\theta\approx 0.17\,{\rm rad}$. We conclude that all inferred $0.05\lesssim\Delta\theta\lesssim 0.2$ values are consistent with the fiducial wobbling jet structure, characterized by a wobbling angle of $\theta_{w}\lesssim 0.3\,{\rm rad}$.
@@ 3.3.1 EP250108a
When no jet afterglow emission is detected, whether due to a large viewing angle, the absence of a relativistic jet, or an intrinsically weak jet, optical signatures that would otherwise be outshone by the jet can be detected. A notable example is EP250108a, which exhibited a fast X-ray transient (Li et al., 2025) followed by a declining optical signal preceding the SN main peak at $t\approx 15\,{\rm days}$ (Eyles-Ferris et al., 2025; Rastinejad et al., 2025; Srinivasaragavan et al., 2025). To explain the optical emission prior to the $^{56}{\rm Ni}$ emission, several models have been proposed, including SN ejecta-CSM interaction and cocoon cooling emission (Eyles-Ferris et al., 2025; Srinivasaragavan et al., 2025; Zhu et al., 2025). However, as we show below, neither model can sufficiently account for the observed week-long decay.
@@
We conclude that in EP250108a, the observed data is inconsistent with the cocoon cooling emission, which decays too fast, favoring disk ejecta as the dominant emission source at $t\gtrsim 3\,{\rm days}$. Collapsar accretion disks (e.g., Zenati et al., 2020; Dean & Fernández, 2024b; Fujibayashi et al., 2024; Menegazzi et al., 2024) and PM outflows (e.g., Suwa & Tominaga, 2015; Wang et al., 2016; Moriya et al., 2017) have been proposed as primary production sites of $^{56}{\rm Ni}$ in stripped-envelope SNe, where much larger amounts are found compared to ordinary core-collapse SNe. The launching of massive disk ejecta, potentially preceded by PM outflows, enhances mixing efficiency and can account for the substantial $^{56}{\rm Ni}$ mass in the outer layers of the ejecta in EP250108a (Rastinejad et al., 2025), as well as in stripped-envelope SNe more broadly (e.g., Maeda & Nomoto, 2003; Maeda et al., 2003; Ashall et al., 2019; Moriya et al., 2020).
Figure 5: (a): Analytic (dashed gray lines) and numerical (solid lines) calculations of cocoon (blue), disk ejecta (green), and total (black lines) cooling emission are compared with the estimated bolometric luminosity of EP250108a (black dots) (Srinivasaragavan et al., 2025) pre-SN peak at $t\approx 15\,{\rm days}$. The temporal evolution indicates that cocoon emission may contribute to the first data point at $t\approx 1\,{\rm day}$, but it decays too rapidly to account for the later observations. The disk ejecta emission dominates the observations at $t\gtrsim 3\,{\rm days}$. (b): $\nu L_{\nu}$ light curve of EP240414a across optical/infrared bands (Srivastav et al., 2025; Sun et al., 2025; van Dalen et al., 2025) presents three phases of optical evolution: cooling emission produces the shallow power-law decline during the first day (blue-shaded background); a subsequent peak powered by a non-thermal afterglow from a wobbling top-hat jet (green-shaded background); and the emergence of thermal emission generated by $^{56}{\rm Ni}-$decay (red-shaded background).

### v2

@@ 1 Introduction
One natural candidate for producing the sub- and mildly relativistic hot plasma is the “cocoon,” generated during the jet’s propagation within the collapsing star (Mészáros & Rees, 2001; Ramirez-Ruiz et al., 2002; Matzner, 2003; Lazzati & Begelman, 2005). Cocoon emission has been predicted to produce vigorous UV/optical signals (Nakar & Piran, 2017), potentially associated with SNe Ib/c (Piran et al., 2019), and is hypothesized as a possible source of fast optical transients (Gottlieb et al., 2022d). Indeed, cocoon emission is a popular interpretation for several puzzling EP transients, including EP240414a (Hamidani et al., 2025b; van Dalen et al., 2025; Zheng et al., 2025), EP240801a (Jiang et al., 2025), EP241021a (Gianfagna et al., 2025), and EP250108a (Eyles-Ferris et al., 2025; Yadav et al., 2025) [^1] Although Yadav et al. (2025) argue against quasi-spherical mildly relativistic cocoons, we note that mildly relativistic cocoons must inherently exhibit some degree of collimation. In fact, the cocoon structure naturally arising during jet propagation constitutes the “structured jet” proposed by Jiang et al. (2025); Yadav et al. (2025) (see Gottlieb et al., 2021, for further discussion on the cocoon-jet interplay)..
@@ 2 Physics of Collapsars
The collapse of rapidly rotating stars leads to the formation of a spinning PM (e.g., Raynaud et al., 2020; Masada et al., 2022; White et al., 2022), which may eventually collapse into a spinning BH. Such a BH can launch relativistic jets via the Blandford-Znajek (BZ; Blandford & Znajek, 1977) mechanism. Sustained jet launching capable of powering GRBs through the BZ mechanism (Popham et al., 1999; MacFadyen & Woosley, 1999) typically requires three conditions (e.g., Symbalisty, 1984; Metzger et al., 2008; Gottlieb et al., 2022a): (i) a dimensionless BH spin parameter $a\gtrsim 0.3$ (Gottlieb et al., 2023a); (ii) the presence of a steady accretion disk surrounding the BH; and (iii) sufficient poloidal magnetic flux available to the BH. Gottlieb et al. (2024) demonstrated that an accretion disk may form around the PM before its collapse; such a disk can inhibit magnetic reconnection, enabling the BH to inherit the PM’s magnetic flux and thus relaxing the third requirement. Consequently, an accretion disk around a moderately spinning BH emerges as the only prerequisite for BZ jet launching. However, the angular momentum profiles of stellar progenitors and the likelihood of disk formation during the PM phase remain uncertain (see e.g., Fryer et al., 2025).
@@
Structure: The collapse results in the formation of a proto-magnetar, which launches subrelativistic outflows (violet) prior to collapsing into a black hole. The ensuing formation of a black hole surrounded by an accretion disk drives moderately collimated, magnetized disk ejecta (turquoise) and a wobbling ultra-relativistic episodic jet powered by the Blandford-Znajek mechanism (yellow). As the jet propagates, it shocks the surrounding stellar material, generating a collimated, mildly relativistic cocoon (blue). An additional subrelativistic SN shockwave from core bounce (magenta) is possible; however, the proto-magnetar and/or disk ejecta are consistent with the energy and velocity distributions observed in stripped-envelope SNe, rendering the SN shockwave unnecessary to account for observations. We note that the illustrated collimation of the cocoon and disk ejecta represents characteristic angles where the bulk of the energy is concentrated. In practice, the outflows mix in a spherical morphology, with the energy distributed as a function of angle.
@@ 2.1.3 Wobbling Jets and Cocoon
If the BH possesses a moderate spin of $a\gtrsim 0.2$, a sufficient amount of magnetic flux will activate the BZ mechanism, launching bipolar relativistic jets. The jets interact with the dense stellar core, producing a shocked cocoon composed of mildly relativistic jet plasma and spherical non-relativistic shocked stellar material (Bromberg et al., 2011). As the cocoon expands, it envelops and interacts with the accretion disk. Large-scale, long-duration, high-resolution 3D general-relativistic magnetohydrodynamic collapsar simulations by Gottlieb et al. (2022b) followed the post-BH formation phase, and demonstrated the self-consistent formation of both the accretion disk and jets from the collapsing star and through the BZ mechanism. These simulations revealed that jet–disk interactions transfer stochastic momentum to the disk, destabilizing it and continuously altering its orientation. Consequently, the jet, which is anchored to the disk’s angular momentum, develops a wobbling motion at its base. [^2] Although only a handful of global collapsar simulations exist due to their high computational cost, we have consistently observed wobbling in all of our global collapsar runs, including direct-horizon collapsars (Gottlieb et al., 2025), with additional setups in preparation. The underlying driver of this behavior is the inevitably strong disk–cocoon interaction. By contrast, in post–compact-object mergers, the disk–cocoon interaction is considerably weaker because there is no sustained infalling envelope as in collapsars, leading to correspondingly weaker wobbling (e.g., Gottlieb et al., 2022c, 2023b). This emerging trend aligns with the expected dependence of wobbling on the strength of the disk–cocoon interaction. Nevertheless, further numerical work is required to establish that this behavior is robust to currently missing physics, such as neutrino-cooled disks in global simulations. Figure 2, taken from Gottlieb et al. (2022b)’s 3D simulations, demonstrates that this motion is imprinted onto the jet’s larger-scale structure and remains observable even after the jet emerges from the collapsing star on a breakout time of a few seconds.
@@ 2.2 Outflow Structure
The relative contribution between the jet-cocoon structure and disk ejecta depends on the BH spin. Provided sufficient magnetic flux is supplied to the BH, either by the PM (Gottlieb et al., 2024) or through dynamo processes (Shibata et al., 2025), two regimes are relevant: [^3] The jet power depends on the BH spin and the magnetic flux. Since collapsar jet launching requires a high magnetic flux (Issa et al., 2025), this imposes a constraint on the BH spin, as high spins would produce jets with luminosities exceeding those observed in GRBs. As a result, spins of $a\gtrsim 0.5$ are ruled out (Gottlieb et al., 2023a). moderately spinning BHs ($0.2\lesssim a\lesssim 0.5$) that launch GRB jets; and slowly spinning BHs ($a\lesssim 0.2$) that do not produce jets capable of powering GRBs. At $a\approx 0.25$, the power of the disk ejecta is comparable to that of the BZ jets, $L_{\rm disk}\approx L_{\rm BZ}$ (McKinney et al., 2012; Narayan et al., 2012; Tchekhovskoy & McKinney, 2012; Bopp & Gottlieb, 2025). Given that jets require a persistent accretion disk, the disk and jet launching durations are also expected to be comparable, leading to similar total energies, $E_{\rm disk}\approx E_{\rm BZ}$. Since jets deposit about half of their energy into the cocoon (Bromberg et al., 2011; Gottlieb et al., 2021), we expect $E_{\rm disk}\approx 2E_{\rm cocoon}\approx 2E_{\rm jet}$. Conversely, in the case of slowly spinning BHs, the disk ejecta becomes the dominant energy channel, $E_{\rm disk}\gg E_{\rm jet}$.
@@
Ultimately, all outflow components interact and merge into a single structured outflow with both radial and angular profiles. Figure 3(a) presents the radially integrated isotropic-equivalent energy as a function of angular separation from the stellar rotation axis. The resulting global angular structure features a flat core with power-law wings (Gottlieb et al., 2021), and remains unaffected by jet wobbling (Gottlieb et al., 2022b), as it reflects radial and azimuthal integrations over multiple jet episodes launched at varying angles. However, the angular structure of individual jet episodes has not been investigated until now. Figure 3(b,c) depicts $E_{\rm iso}$ (b) and Lorentz factor (c) across a shell containing a single bipolar wobbling jet episode. Both jets show a prominent deviation from the rotational axis due to the wobbling motion. In stark contrast to the smooth core+power-law structure seen in the global outflow structure (dashed lines in panel b), single jet episodes exhibit a distinct core and steep power-law wings, whose power-law indices exceed $\sim 5$ for both the energy and Lorentz factor. Accordingly, the angular profile of single jet episodes is more characteristic of a top-hat jet than a structured jet. We note that the relatively low jet Lorentz factor of $\Gamma\lesssim 10$ in Fig. 3(c) arises from the floor values imposed in numerical simulations, which constrain the asymptotic Lorentz factor. Nevetheless, Gottlieb et al. (2022b) showed that a higher asymptotic Lorentz factor, as suggested by observations, does not alter the global outflow morphology, and is expected to sharpen the episodic top-hat structure of the relativistic jet relative to the mildly relativistic cocoon, thereby further accentuating its top-hat profile.
@@ 3.2.1 Challenges of Proposed Models
The detection of GRBs, along with the direct association of some events with SNe Ic-BL, supports a collapsar origin and thus motivates a focus on collapsar models rather than other events such as accreting white dwarfs or binaries (Wu et al., 2025), or tidal disruption events (Shu et al., 2025), whose afterglows evolve over considerably longer timescales (e.g., Alexander et al., 2020). As density fluctuations in the CSM cannot account for the observed sharp features (Nakar & Granot, 2007), a few alternative models have been proposed to explain the afterglow rebrightening. Two widely discussed models proposed to explain the rebrightening are off-axis emission from a jet with angular structure (e.g., Yadav et al., 2025; Zheng et al., 2025), and refreshed shocks (e.g., Busmann et al., 2025; Ricci et al., 2025; Srivastav et al., 2025), or a combination of both (Gianfagna et al., 2025). However, both of these models face significant challenges.
Structured jets can naturally explain the absence of $\gamma$-ray emission when the observer’s line of sight initially misses the jet core, and the rebrightening emerges as the jet core gradually enters the line of sight. However, the very steep rise observed in EP241021a ($L_{\nu}\sim t^{4}$; Busmann et al., 2025) and EP240414a ($L_{\nu}\sim t^{3}$; Zheng et al., 2025), cannot be reproduced by a smooth angular jet structure (Lamb et al., 2021), unless the jet is composed of discrete segments (Gianfagna et al., 2025) with distinct Lorentz factors (Shu et al., 2025), similar to top-hat jets (Granot et al., 2002; Nakar et al., 2002). Such a segmented configuration of two-component relativistic jets (see Granot, 2005; Beniamini et al., 2020; Li et al., 2023), previously proposed to explain rebrightening features (Liang et al., 2013), naturally arises from wobbling top-hat jets, as we demonstrated in §2.2, and as we argue in §3.2.2, likely accounts for the observed sharp rebrightening behavior. In contrast, the assumed structures in Gianfagna et al. (2025); Yadav et al. (2025); Zheng et al. (2025) of jets with only an angular profile formed through continuous interaction with the ambient medium, will produce a monotonic jet+cocoon structure (see jet structure profiles in Gottlieb et al., 2021), leading to a gradual and prolonged rise, similar to that in GW170817 (e.g., Mooley et al., 2018). Similar challenges arise for radial structures, which are characteristic of choked jets (see §3.3.2).
@@ 3.2.2 Wobbling Jets
The high-redshift ($z=4.9$) event EP240315a exhibited a rebrightening at $t_{\rm rb}\approx 13\,{\rm hr}$. Assuming the kinetic energy is comparable to the high-energy emission, $E_{k,\rm iso}\approx E_{\rm iso}\approx 10^{54}\,\,{{\rm erg}}$, and $n=1\,\,{{\rm cm}}^{-3}$, we find from Equation (1) that $\Gamma\approx 20$, implying the observer was located $\Delta\theta\approx 0.05\,{\rm rad}$ off-axis from the most energetic segment of the wobbling jet. The jet afterglow in EP240414a at $z=0.4$ indicates $E_{k,\rm iso}\approx 10^{51}\,\,{{\rm erg}}$ with $n=1\,\,{{\rm cm}}^{-3}$ for the initially observed decelerating jet prior to brightening (Sun et al., 2025). Assuming a typical total (over all episodes) jet isotropic equivalent energy of $E_{k,\rm iso}\approx 10^{53}\,\,{{\rm erg}}$ at $t_{\rm rb}\approx 1\,{\rm day}$, we find $\Gamma\approx 7$, implying that $\Delta\theta\approx 0.15\,{\rm rad}$. For EP241021a ($z=0.75$), which exhibited a rebrightening at $t_{\rm rb}=6\,{\rm days}$, we adopt $E_{k,\rm iso}=10^{53}\,\,{{\rm erg}}$ and $n=10^{-3}\,\,{{\rm cm}}^{-3}$ based on afterglow modeling by Yadav et al. (2025). This yields $\Gamma\approx 9$, corresponding to $\Delta\theta\approx 0.1\,{\rm rad}$. A secondary, weaker rebrightening, which could be powered by a weaker wobbling jet episode, was observed at $t_{\rm rb}=20.7\,{\rm days}$ (Shu et al., 2025), which corresponds to $\Delta\theta\approx 0.17\,{\rm rad}$. We conclude that all inferred $0.05\lesssim\Delta\theta\lesssim 0.2$ values are consistent with the fiducial wobbling jet structure, characterized by a wobbling angle of $\theta_{w}\lesssim 0.3\,{\rm rad}$.
@@ 3.3.1 EP250108a
When no jet afterglow emission is detected, whether due to a large viewing angle, the absence of a relativistic jet, or an intrinsically weak jet, optical signatures that would otherwise be outshone by the jet can be detected. A notable example is EP250108a, which exhibited a fast X-ray transient (Li et al., 2025) followed by a declining optical signal preceding the SN main peak at $t\approx 12\,{\rm days}$ (Eyles-Ferris et al., 2025; Rastinejad et al., 2025; Srinivasaragavan et al., 2025). To explain the optical emission prior to the $^{56}{\rm Ni}$ emission, several models have been proposed, including SN ejecta-CSM interaction and cocoon cooling emission (Eyles-Ferris et al., 2025; Srinivasaragavan et al., 2025; Zhu et al., 2025). However, as we show below, neither model can sufficiently account for the observed week-long decay.
@@
We conclude that in EP250108a, the observed data are inconsistent with the cocoon cooling emission, which decays too fast, favoring disk ejecta as the dominant emission source during $3\,{\rm days}\lesssim t\lesssim 7\,{\rm days}$, prior to the emergence of the main SN signal. While the simulation in Bopp & Gottlieb (2025) does not track the gas composition of the disk ejecta, axisymmetric studies have shown that collapsar accretion disks (e.g., Zenati et al., 2020; Dean & Fernández, 2024b; Fujibayashi et al., 2024; Menegazzi et al., 2024) and PM outflows (e.g., Suwa & Tominaga, 2015; Wang et al., 2016; Moriya et al., 2017) can serve as primary production sites of $^{56}{\rm Ni}$ in stripped-envelope SNe, yielding significantly larger amounts than those found in ordinary core-collapse SNe. The launching of massive disk ejecta, potentially preceded by PM outflows, enhances mixing efficiency and can account for the substantial $^{56}{\rm Ni}$ mass in the outer layers of the ejecta in EP250108a (Rastinejad et al., 2025), as well as in stripped-envelope SNe more broadly (e.g., Maeda & Nomoto, 2003; Maeda et al., 2003; Ashall et al., 2019; Moriya et al., 2020).
Figure 5: (a): Analytic (dashed gray lines) and numerical (solid lines) calculations of cocoon (blue), disk ejecta (green), and total (black lines) cooling emission are compared with the estimated bolometric luminosity of EP250108a (black dots) (Srinivasaragavan et al., 2025). The temporal evolution indicates that cocoon emission may contribute to the first data point at $t\approx 1\,{\rm day}$, but it decays too rapidly to account for the later observations. The disk ejecta emission dominates the observations at $3\,{\rm days}\lesssim t\lesssim 7\,{\rm days}$, before the main SN peak at $t\approx 12\,{\rm days}$. (b): $\nu L_{\nu}$ light curve of EP240414a across optical/infrared bands (Srivastav et al., 2025; Sun et al., 2025; van Dalen et al., 2025) presents three phases of optical evolution: cooling emission produces the shallow power-law decline during the first day (blue-shaded background); a subsequent peak powered by a non-thermal afterglow from a wobbling top-hat jet (green-shaded background); and the emergence of thermal emission generated by $^{56}{\rm Ni}-$decay (red-shaded background).